Two tiny spheres carrying charges of 1 mC and 3 mC are placed 8cm apart in air.The electrostatic potential at the midpoint of the line joining the two charges?
OPTIONS:
9×10²V
9×10³V
9×10⁴V
9×10⁵V
9×10⁶V

Grade:12th Pass

1 Answers

Eshan
5 years ago
Dear student,

Distance of the mid point from each sphere=4cm=0.04m

Potential at mid point due to 1mC charge=\dfrac{(9\times 10^9)(0.001)}{0.04}V

Potential at mid point due to 3mC charge=\dfrac{(9\times 10^9)(0.003)}{0.04}V

Hence total potential at the point=9\times 10^8V